# 22footnotetext: Corresponding author

### v1

@@ 3.5.5. Miscellaneous
Beyond the primary tasks discussed above, Mamba-based models have been employed in a range of other medical imaging applications. Table 7 outlines models and their parameters along with their descriptions and code availability. Notable mamba-based research papers in this field include:
@@
Xie et al. (Xie et al., 2024b) introduced Prompt-Mamba for polyp segmentation in colonoscopy images. Polyp segmentation is crucial for early cancer detection but it is challenging due to variations in size, shape, and color. Existing models struggle with irregular shapes, small samples and generalization to unseen data. Prompt-Mamba addresses these limitations by combining Vision-Mamba which is known for its robust image feature extraction with prompt technology for improved generalization. The lightweight architecture includes an image encoder with ViM layers, a prompt encoder and a mask decoder. It also uses a combination of Focalloss and Diceloss functions to effectively handle imbalanced data and measure segmentation quality. Prompt-Mamba surpasses existing state-of-the-art methods by an average of 5% across six datasets. It demonstrates exceptional performance on both validation and unseen datasets.
Table 7. Overview of Miscellaneous Models
@@
P-Mamba (Ye et al., 2024) 52.77 ViM Pediatric cardiac imaging
Prompt-Mamba (Xie et al., 2024b) 102 ViM Polyp analysis with prompt technologies
ClinicalMamba (Yang et al., 2024c) - Mamba Language model with prompt based fine-tuning
Vivim (Yang et al., 2024e) - ST-Mamba Video analysis of multiple tumor
VM-DDPM (Ju and Zhou, 2024) 15.2 VMamba Synthesis of MRI-X ray images
MD-Dose (Fu et al., 2024) - Mamba Predicts dosage distribution for tumor patients
@@
camera tracker (Zhang et al., 2024c)
@@
Yang et al. (Yang et al., 2024c) proposed ClinicalMamba, a specialized language model based on Mamba. ClinicalMamba is pre-trained on longitudinal clinical notes from MIMIC-III to achieve notable speed and performance benchmarks. It surpasses established clinical language models as well as large language models such as GPT-4 in longitudinal clinical tasks. The researchers pre-trained ClinicalMamba using a causal language modeling objective on a dataset of de-identified clinical notes from MIMIC-III. This dataset encompassed 82,178 hospital visits with all associated notes aggregated longitudinally. The model leverages a selective SSM with a context window of 16k tokens, effectively covering over 98% of the visits within the dataset. By specializing in medical domain and pre-training on longitudinal data, ClinicalMamba captures unique characteristics of clinical narratives. The paper highlights potential applications in tasks such as cohort selection for clinical trials and ICD coding, where the model can interpret clinical narratives and extract relevant information. Ye et al. (Ye et al., 2024) devised P-Mamba, a novel deep learning model to address pediatric echocardiographic left ventricular segmentation. It tackles two key issues: computational efficiency and noise interference. The model employs a dual-branch architecture. ViM encoder branch focuses on efficiency by capturing global dependencies in the image. DWT-based PMD encoder branch, leveraging a technique originally used for image de-noising (DWT-based PMD) block, specifically targets noise suppression while preserving local features crucial for segmentation. This branch utilizes an anisotropic diffusion equation to achieve this balance. Finally, decoders within the model (SegHead and FCNHead) upsample features from both branches to generate segmentation masks.
Fu et al. (Fu et al., 2024) introduced MD-Dose for predicting radiation therapy dose distributions using a Mamba-based diffusion model. MD-Dose applies Mamba within diffusion model for both denoising and encoding tasks. The forward process of MD-Dose involves adding Gaussian noise to dose distribution maps, while the reverse process uses a noise predictor to reconstruct the original maps from noise. Mamba encoder in MD-Dose is used to extract structural information from CT images and then integrating it with noise prediction to enhance the localization of dose regions in Planning Target Volume (PTV) and Organs At Risk (OAR). The experiments conducted on a dataset of 300 thoracic tumor patients demonstrate that MD-Dose outperforms existing models in both accuracy and inference speed. The efficiency of architecture highlights its reduced parameter count and faster inference time compared to other methods which makes it a significant advancement in automated radiotherapy planning.
Yang et al. (Yang et al., 2024e) proposed Vivim, a novel deep learning framework shown in Figure 29 for medical video object segmentation. Vivim addresses challenges of capturing spatiotemporal information and handling long-range dependencies in medical videos. The core of Vivim lies in its Video Vision Mamba (ST-Mamba) module, which leverages a transformer architecture with a selective scan mechanism inspired by SSMs. ST-Mamba efficiently captures long-range temporal dependencies within video sequences. The hierarchical encoder utilizes temporal Mamba blocks to extract multi-scale feature sequences, while a lightweight CNN-based decoder head fuses these features and predicts segmentation masks. Vivim offers several advantages. Firstly, it integrates SSMs into medical video object segmentation for the first time, achieving superior performance and speed compared to existing methods. Secondly, the boundary-aware constraint improves the model’s ability to discriminate between ambiguous tissues, a frequent challenge in medical videos. The paper also introduces a new benchmark dataset VTUS which includes annotated thyroid ultrasound videos. Extensive evaluations demonstrate that Vivim outperforms state-of-the-art segmentation models on both VTUS and polyp colonoscopy datasets. It highlights its effectiveness in medical video object segmentation.
Figure 29. Architecture of Vivim (Yang et al., 2024e)
@@
Ju et al. (Ju and Zhou, 2024) proposed VM-DDPM, a U-Net based architecture which performs forward noising and reverse denoising process as proposed in Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020). The authors introduced SSLayer in this architecture which contains residual block and modification of VSS block called Multi-level State Space (MSS) block. MSS block incorporates time step embedding for diffusion process and Cross-Scan Module (CSM) which is a modification of S6 scan. To adapt the model for synthesis of new images, the image patches are randomly shuffled and then scanned, similar to SS2D in VSS. This scanning process requires zero parameters. Zhang et al. (Zhang et al., 2024c) introduced a low-cost dual-camera tracker for endoscopy skill evaluation in mechanical simulators. The tracker solves problems faced by existing methods such as lack of consistency and high cost. It achieves accurate and reliable 3D endoscope tip position feedback. The core of the system lies in cross-Camera Mutual Template (CMT) strategy that utilizes information from both cameras to maintain tracking consistency. Motion-Guided Prediction Head (MMH) based on Mamba integrates historical motion data with visual tracking. It enhances its performance in scenarios with occlusions and distortions. Furthermore, the tracker includes a Vision-Motion integrator that combines motion information with visual features for improved 3D localization. This integration is done by a mechanism named Multi-KV cross-attention. The paper also introduces Selective Scan SSMs within the tracking which leads to superior performance compared to traditional SSMs. To evaluate the tracker’s effectiveness, researchers developed a customized mechanical gastric simulator with calibrated cameras and an annotated dataset. The combination of CMT, MMH, Vision-Motion, and SSMs demonstrates the tracker’s potential for improving endoscopy training in mechanical simulators.
4. Datasets
@@
Application in medical domain depends on a wide variety of datasets to improve research and innovation in the field of medical imaging, diagnostics and treatment planning. These datasets cover multiple medical specialties and imaging modalities. In this section, we offer a comprehensive overview of some of the most significant datasets utilized in Mamba-based models across various medical fields.
  1. i.
    BraTS2023 Dataset (Menze et al., 2015; et al., 2024b; Bakas et al., 2017) consists of a diverse collection of structural MRI sequences from pediatric patients diagnosed with high-grade glioma, gathered across multiple institutions.
  2. ii.
    AIIB2023 (Nan et al., 2022; Li et al., 2022) establishes a new benchmark for AI-based solutions in airway modeling, specifically targeting lung fibrosis disease. It includes 120 high-resolution CT (HRCT) scans, each with rare expert annotations.
  3. iii.
    CRC-500 Dataset (Xing et al., 2024) is a comprehensive and large-scale resource designed for 3D colorectal cancer segmentation, comprising 500 meticulously annotated 3D Computed Tomography (CT) scans.
  4. iv.
    ISIC2017 (Codella et al., 2018) challenge offers a comprehensive dataset for lesion image analysis, including approximately 2,000 training images to facilitate development across three key components of the task.
  5. v.
    ISIC2018 (Codella et al., 2019; Tschandl et al., 2018) comprises over 12,500 images distributed across three tasks, each designed to address distinct aspects of dermoscopic image analysis.
  6. vi.
    Spleen Dataset (et al., 2022, 2019a) comprises 61 portal-venous phase CT scans from patients undergoing chemotherapy for liver metastases, focusing on the spleen as the target ROI.
  7. vii.
    CVC-ClinicDB Dataset (Bernal et al., 2015) features 612 high-resolution images extracted from 31 colonoscopy sequences, each with a resolution of $384\times 288$ pixels.
  8. viii.
    3D Abdomen CT Dataset (et al., 2023c) comprises of 2,300 CT scans and is part of the MICCAI 2022 FLARE Challenge, focused on segmentation of 13 abdominal organs including pancreas,liver, kidneys, spleen, and more.
  9. ix.
    2D Abdomen MRI Dataset (Ji et al., 2022) also known as AMOS, is a large-scale dataset designed for abdominal organ segmentation, offering 100 MRI scans and 500 CT scans.
  10. x.
    Endoscopy Dataset (et al., 2019b) or ROBUST-MIS, originates from Robotic Instrument Segmentation Sub-Challenge of the MICCAI 2017 EndoVis Challenge.
  11. xi.
    Microscopy Dataset (Ma et al., 2024) was a part of the NeurIPS 2022 Cell Segmentation Challenge, focusing on cell segmentation across various microscopy images sourced from over 20 biology laboratories and covering more than 50 distinct biological experiments.
  12. xii.
    ACDC MRI Cardiac Dataset (Bernard et al., 2018) comprises 150 exams divided into five groups, including four pathological and one healthy group, with each patient’s data including weight, height, and diastolic and systolic phase information.
  13. xiii.
    Synapse Multi-Organ Abdominal CT Dataset  [^5] Synapse Multi-Organ Abdominal CT Dataset includes 50 abdominal CT scans collected under Institutional Review Board (IRB) supervision from ventral hernia and colorectal cancer studies during the portal venous contrast phase, with detailed labeling of thirteen abdominal organs verified by a radiologist.
  14. xiv.
    $\mathbf{PH^{2}}$ Dataset (Mendonça et al., 2013) includes 200 dermoscopic images of melanocytic lesions, with 80 common nevi, 40 melanomas, and 80 atypical nevi acquired at the Dermatology Service of Hospital Pedro Hispano in Portugal.
  15. xv.
    The LiTS Dataset (et al., 2023d), organized in conjunction with IEEE ISBI 2017 and MICCAI 2017-2018, features a diverse image dataset that includes both primary and secondary liver tumors of varying sizes and appearances.
  16. xvi.
    Montgomery and Shenzhen Dataset (Jaeger et al., 2014) consists of 138 frontal chest X-rays from Montgomery County, including 80 normal and 58 tuberculosis (TB) cases. Shenzhen dataset, compiled with Shenzhen No.3 People’s Hospital and Guangdong Medical College, includes 336 cases with TB manifestations and 326 normal cases.
  17. xvii.
    The Brain MRI Multiple Sclerosis Dataset (M. Muslim e al., 2022) of 2D FLAIR MRI has scans from 60 patients represents a valuable resource for studying Multiple Sclerosis (MS) in a clinical setting. CMR×Recon MICCAI-2023 (et al., 2023a) challenge data has cardiac Magnetic Resonance Imaging (MRI) data from 300 subjects, widely used for both reconstruction and segmentation tasks.
  18. xviii.
    Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Jack et al., 2008; Lian et al., 2020) is an ongoing, multisite observational study involving healthy older adults, individuals with mild cognitive impairment (MCI), and those with Alzheimer’s disease.
  19. xix.
    Otoscopy Dataset (Zeng et al., 2021) comprises images from 41,056 patients diagnosed in the Department of Otolaryngology at the People’s Hospital of Shenzhen Baoan District, captured with standard endoscopes and digital endovision camera systems.
  20. xx.
    PathMNIST (Yang et al., 2023, 2021) consists of 100,000 non-overlapping image patches from hematoxylin and eosin-stained histological images and an additional test set of 7,180 patches from a clinical center for predicting survival outcomes in colorectal cancer.
  21. xxi.
    CAMELYON16 (Bandi, [n. d.]; Bejnordi et al., 2017) includes 270 Whole-Slide Images (WSIs) for training (159 normal and 111 with tumors) and 129 WSIs for testing, aimed at detecting metastases in lymph node sections using eosin (H&E) staining and hematoxylin.
  22. xxii.
    The Colorectal Cancer Histopathology Dataset (Loughrey et al., 2020) developed by International Collaboration on Cancer Reporting (ICCR), provides pathology reports on colorectal cancer surgical specimens.
  23. xxiii.
    SynthRAD (Guo et al., 2024) evaluates methods across different MRI sequences and acquisition settings with 60 MRI, 20 CT and 10 CBCT scans for train, test, and validation respectively, along with 540 paired MRI-CT and 540 CBCT-CT sets, and includes a mask for evaluation with target information for validation and test sets kept confidential.
  24. xxiv.
    The FastMRI Dataset (Zbontar et al., 2018), consisting of 584 3D knee MRI scans, was utilized for the experiments. To ensure data quality, 20 central coronal slices were selected from each case and cropped to 320$\times$320 pixels. The train, test and validation split is in 7:2:1 ratio. The provided emulated single-coil data served as the ground truth for single-coil complex-value reconstruction.
  25. xxv.
    Low-Dose CT Image and Projection Dataset (Moen et al., 2021) were divided into chest and abdomen subsets. The chest subset consisted of non-contrast scans for pulmonary nodule screening, while the abdomen subset included contrast-enhanced scans for metastatic liver lesion detection. Each subset contained 40 cases, split into training and testing sets. Sparse-view sinograms were created using torch-radon in a fan-beam CT geometry with 736 detectors and 60 projection views. The reconstructed image resolution was 512$\times$512 pixels.
@@
Table 8. Overview of Datasets, Tasks Performed, and Targeted Areas in Medical Imaging and Analysis
Dataset Task Performed Targeted Area
BraTS2023 (Menze et al., 2015; et al., 2024b; Bakas et al., 2017) Segmentation Brain
AIIB2023 (Nan et al., 2022; Li et al., 2022) Segmentation Lungs
CRC-500 Dataset (Xing et al., 2024) Segmentation Colon, Rectum
ISIC2017 (Codella et al., 2018) Segmentation Skin
ISIC2018 (Codella et al., 2019; Tschandl et al., 2018) Segmentation Skin
MICCAI-2023 (et al., 2024c) Segmentation Heart
Spleen (et al., 2022, 2019a) Segmentation Spleen
CVC-ClinicDB (Bernal et al., 2015) Segmentation Intestine
3D Abdomen CT dataset (et al., 2023c) Segmentation Liver, Kidney, Spleen, etc.
2D Abdomen MRI Dataset (Ji et al., 2022) Segmentation Kidney, Gallbladder, Esophagus,
Liver, Pancreas, Adrenal gland, etc.
@@
Endoscopy Dataset (et al., 2019b) Segmentation Surgical Instruments
@@
Microscopy Dataset (Ma et al., 2024) Segmentation Cell
@@
ACDC MRI Cardiac Dataset (Bernard et al., 2018) Segmentation Heart
@@
Synapse Multi-Organ Abdominal Segmentation Pancreas, Spleen , Liver
CT Dataset 55footnotemark: 5
$PH^{2}$ Dataset (Mendonça et al., 2013) Segmentation Skin
LiTS Dataset (et al., 2023d) Segmentation Liver
@@
Montgomery and Shenzhen Segmentation Lungs
Dataset(Jaeger et al., 2014)
Brain MRI Multiple Sclerosis Segmentation Brain
Dataset(M. Muslim e al., 2022)
Alzheimer’s Disease Neuroimaging Classification Brain
(ADNI) Dataset (Jack et al., 2008; Lian et al., 2020)
Otoscopy (Zeng et al., 2021) Classification Ear
PathMNIST (Yang et al., 2023, 2021) Classification Colon Pathology
@@
Camelyon16 (Bandi, [n. d.]; Bejnordi et al., 2017) Classification Lymph node
@@
Colorectal Cancer Histopathology Classification Colon, Rectum
Dataset (Loughrey et al., 2020)
SR-Reg (SynthRAD Registration) Registration Brain
dataset (Guo et al., 2024)
FastMRI Knee Dataset (Zbontar et al., 2018) Restoration/ Reconstruction Knee
Low-Dose CT Image and Restoration/ Reconstruction Head, Chest, Abdomen
Projection Datasets (Moen et al., 2021)
Table 8 provides a detailed summary of various datasets used in medical imaging, specifying tasks performed and targeted areas, which includes both anatomical organs and surgical instruments. It highlights diversity of dataset in medical applications, offering insight into how different challenges in medical diagnostics and treatment are addressed through targeted imaging and analysis.
5. Experiments and Results

### v3

@@ 3.5.5 Miscellaneous
Beyond primary tasks like segmentation, classification, and image restoration discussed in earlier sections, Mamba-based models have also been applied to a wide range of medical applications. These tasks include video analysis, medical language understanding, and image generation using diffusion models. In all these cases, Mamba helps to improve efficiency and better capture long-range relationships in data. Table 7 outlines models and their parameters along with their descriptions and code availability. Notable mamba-based research papers in this field include
@@
Xie et al. [69] introduced Prompt-Mamba for polyp segmentation in colonoscopy images. Polyp segmentation is crucial for early cancer detection but it is challenging due to variations in size, shape, and color. Prompt-Mamba addresses existing limitations such as various sample sizes and generalization to unseen data by combining Vision-Mamba which act as an image feature extractor with prompt technology for improved generalization. The lightweight architecture includes an image encoder with ViM layers, a prompt encoder and a mask decoder. It uses combination of Focal and Dice loss to handle class imbalance and measure segmentation quality.
Table 7: Overview of Miscellaneous Models
@@
P-Mamba [71] 52.77 ViM Pediatric cardiac imaging
Prompt-Mamba [69] 102 ViM Polyp analysis with prompt technologies
ClinicalMamba [70] - Mamba Language model with prompt based fine-tuning
Vivim [73] - ST-Mamba Video analysis of multiple tumor
VM-DDPM [74] 15.2 VMamba Synthesis of MRI-X ray images
MD-Dose [72] - Mamba Predicts dosage distribution for tumor patients
@@
camera tracker [143]
Yang et al. [70] proposed ClinicalMamba, a specialized language model based on Mamba. ClinicalMamba is pre-trained on longitudinal clinical notes from MIMIC-III to achieve notable speed and performance benchmarks. It surpasses established clinical language models as well as large language models such as GPT-4 in longitudinal clinical tasks. The researchers pre-trained ClinicalMamba using a causal language modeling objective on a dataset of de-identified clinical notes from MIMIC-III. By specializing in medical domain and pre-training on longitudinal data, ClinicalMamba captures unique characteristics of clinical narratives.
@@
Ye et al. [71] devised P-Mamba, a novel deep learning model to address pediatric echocardiographic left ventricular segmentation. It tackles two key issues: computational efficiency and noise interference. The model employs a dual-branch architecture. ViM encoder branch focuses on efficiency by capturing global dependencies in the image. DWT-based PMD encoder branch, leveraging a technique originally used for image de-noising (DWT-based PMD) block, specifically targets noise suppression while preserving local features crucial for segmentation. This branch utilizes an anisotropic diffusion equation to achieve this balance. Finally, decoders within the model (SegHead and FCNHead) upsample features from both branches to generate segmentation masks. Fu et al. [72] introduced MD-Dose for predicting radiation therapy dose distributions using a Mamba-based diffusion model. MD-Dose applies Mamba within diffusion model for both denoising and encoding tasks. The forward process of MD-Dose involves adding Gaussian noise to dose distribution maps, while the reverse process uses a noise predictor to reconstruct the original maps from noise. Mamba encoder in MD-Dose is used to extract structural information from CT images and then integrating it with noise prediction to enhance the localization of dose regions in Planning Target Volume (PTV) and Organs At Risk (OAR).
@@
Yang et al. [73] proposed Vivim, a novel deep learning framework for medical video object segmentation. Vivim addresses challenges of capturing spatiotemporal information and handling long-range dependencies in medical videos. The core of Vivim lies in its Video Vision Mamba (ST-Mamba) module, which leverages a transformer architecture with a selective scan mechanism inspired by SSMs. ST-Mamba efficiently captures long-range temporal dependencies within video sequences. The hierarchical encoder utilizes temporal Mamba blocks to extract multi-scale feature sequences, while a lightweight CNN-based decoder head fuses these features and predicts segmentation masks. The paper also introduces a new benchmark dataset VTUS which includes annotated thyroid ultrasound videos. Ju et al. [74] proposed VM-DDPM, a U-Net based architecture which performs forward noising and reverse denoising process as proposed in Denoising Diffusion Probabilistic Models (DDPM) [162]. The authors introduced SSLayer in this architecture which contains residual block and modification of VSS block called Multi-level State Space (MSS) block. MSS block incorporates time step embedding for diffusion process and Cross-Scan Module (CSM) which is a modification of S6 scan. Zhang et al. [143] introduced a low-cost dual-camera tracker for endoscopy skill evaluation in mechanical simulators. The tracker solves problems faced by existing methods such as lack of consistency and high cost. It achieves accurate and reliable 3D endoscope tip position feedback. The core of the system lies in cross-Camera Mutual Template (CMT) strategy that utilizes information from both cameras to maintain tracking consistency. Motion-Guided Prediction Head (MMH) based on Mamba integrates historical motion data with visual tracking. Furthermore, the tracker includes a Vision-Motion integrator that combines motion information with visual features for improved 3D localization. This integration is done by a mechanism named Multi-KV cross-attention.
4 Datasets
@@
Application in medical domain depends on a wide variety of datasets to improve research and innovation in the field of medical imaging, diagnostics and treatment planning. These datasets cover multiple medical specialties and imaging modalities. In this section, we offer a comprehensive overview of some of the most significant datasets utilized in Mamba-based models across various medical fields. Table 8 provides a detailed summary of various datasets used in medical imaging, specifying tasks performed and targeted areas, which includes both anatomical organs and surgical instruments. It highlights diversity of dataset in medical applications, offering insight into how different challenges in medical diagnostics and treatment are addressed through targeted imaging and analysis.
@@
Table 8: Overview of Datasets, Tasks Performed, and Targeted Areas in Medical Imaging and Analysis
Dataset Task Performed Targeted Area Description
BraTS2023 [41, 42, 43] Segmentation Brain Collection of multi-parametric MRI scans
AIIB2023 [44, 45] Segmentation Lungs Consists of 120 high-resolution CT scans
CRC-500 Dataset [16] Segmentation Colon, Rectum 500 meticulously annotated 3D CT scans
ISIC2017 [46] Segmentation Skin 2,000 training images across 3 components
ISIC2018 [47, 48] Segmentation Skin 12,500 images distributed across three tasks
MICCAI-2023 [163] Segmentation Heart Contains 120 training, 60 validation and
120 test data
Spleen [164, 165] Segmentation Spleen Includes 61 portal-venous phase CT scans
CVC-ClinicDB [49] Segmentation Intestine Comprises 612 high-resolution images
3D Abdomen CT dataset [50] Segmentation Liver, Kidney, Consists of 2,300 CT scans
Spleen, etc.
2D Abdomen MRI Dataset [51] Segmentation Kidney, Pancreas, Contains 100 MRI scans and 500 CT scans
Gallbladder, Liver,
Adrenal gland,
Esophagus, etc.
@@
Endoscopy Dataset [52] Segmentation Surgical Includes 10,040 annotated images from
Instruments three different surgery types
@@
Microscopy Dataset [53] Segmentation Cell Training set consists of 1000 labeled images,
1712 unlabeled images and 13 unlabeled
WSI and validation set contains 101 images
including 1 WSI
@@
ACDC MRI Cardiac Dataset [54] Segmentation Heart Includes MRI images of 150 patients
@@
Synapse Multi-Organ Abdominal Segmentation Pancreas, Spleen , Comprises 50 abdominal CT scans
CT Dataset 55footnotemark: 5 Liver
$PH^{2}$ Dataset [166] Segmentation Skin Contains 200 dermoscopic images of
melanocytic lesions, 130 CT scans for
training and 70 CT scans for testing
LiTS Dataset [55] Segmentation Liver Consists of 100 training, 10 validation
and 21 test samples
@@
Montgomery and Shenzhen Segmentation Lungs Consists of 138 frontal chest X-rays
Dataset[56]
Brain MRI Multiple Sclerosis Segmentation Brain Includes 60 samples of MRI images
Dataset[57]
Alzheimer’s Disease Neuroimaging Classification Brain Non-converters (NC) and Alzheimer’s
(ADNI) Dataset [58, 59] Disease (AD) classification task
Otoscopy [60] Classification Ear Comprises images from 41,056 patients
PathMNIST [61, 62] Classification Colon Pathology Consists of 100,000 non-overlapping
image patches
@@
Camelyon16 [63, 64] Classification Lymph node Includes 270 WSIs for train (159 normal and
111 with tumors) and 129 WSIs for test
@@
Colorectal Cancer Histopathology Classification Colon, Rectum Developed by International Collaboration
Dataset [65] on Cancer Reporting (ICCR)
SR-Reg (SynthRAD Registration) Registration Brain Includes 60 MRI, 20 CT and 10 CBCT scans
dataset [66] for train, test and validation along with 540
paired MRI-CT and 540 CBCT-CT sets
FastMRI Knee Dataset [67] Restoration/ Knee Contains 584 3D knee MRI scans and dataset
Reconstruction is split in 7:2:1 ratio
Low-Dose CT Image and Restoration/ Head, Chest, Comprises 99 head scans, 100 chest scans
Projection Datasets [68] Reconstruction Abdomen and 100 abdomen scans
5 Experiments and Results